# Prompt Optimization with EASE? Efficient Ordering-aware Automated Selection of Exemplars

### v1

@@ Abstract
Large language models (LLMs) have shown impressive capabilities in real-world applications. The capability of in-context learning (ICL) allows us to adapt an LLM to downstream tasks by including input-label exemplars in the prompt without model fine-tuning. However, the quality of these exemplars in the prompt greatly impacts performance, highlighting the need for an effective automated exemplar selection method. Recent studies have explored retrieval-based approaches to select exemplars tailored to individual test queries, which can be undesirable due to extra test-time computation and an increased risk of data exposure. Moreover, existing methods fail to adequately account for the impact of exemplar ordering on the performance. On the other hand, the impact of the instruction, another essential component in the prompt given to the LLM, is often overlooked in existing exemplar selection methods. To address these challenges, we propose a novel method named EASE, which leverages the hidden embedding from a pre-trained language model to represent ordered sets of exemplars and uses a neural bandit algorithm to optimize the sets of exemplars while accounting for exemplar ordering. Our EASE can efficiently find an ordered set of exemplars that performs well for all test queries from a given task, thereby eliminating test-time computation. Importantly, EASE can be readily extended to jointly optimize both the exemplars and the instruction. Through extensive empirical evaluations (including novel tasks), we demonstrate the superiority of EASE over existing methods, and reveal practical insights about the impact of exemplar selection on ICL, which may be of independent interest. Our code is available at https://github.com/ZhaoxuanWu/EASE-Prompt-Optimization.
@@ 1 Introduction
00footnotetext: * Equal contribution.
Large language models (LLMs) have recently drawn significant attention and have been widely deployed in various real-world scenarios [18, 20, 40] due to their strong capabilities. Of note, a particularly impressive capability of LLMs is in-context learning (ICL): LLMs can learn from a handful of input-label demonstrations (i.e., data exemplars) included in its prompt to perform a downstream task [2, 15]. ICL allows us to adapt an LLM to a downstream task without fine-tuning the model parameters [13, 21]. However, the ICL performance is heavily dependent on the data exemplars in the prompt [1, 15, 29]. Therefore, to maximize the ICL performance, it is of paramount importance to carefully select the set of exemplars. Unfortunately, exemplar selection for ICL is challenging because the mechanism of ICL is complicated and unknown, and this difficulty is further aggravated for black-box LLMs to which we only have API access (e.g., GPT-4 [24], Gemini Ultra [32]).
A large number of existing works on exemplar selection for ICL have considered retrieval-based methods [1, 15, 29]. Specifically, they aim to develop a retriever model such that for every test query, the retriever model retrieves the corresponding best set of exemplars tailored to this particular query [1, 38]. Therefore, at test time, the retriever needs to be deployed for every test query, which could incur significant additional computational overheads especially when the number of test queries is large [17]. This is likely to hinder their ease of adoption in practice because of the prolonged response time and the additional complexity from running the retriever for every query. Moreover, another drawback of these retrieval-based methods is that they may lead to increased privacy risks. Compared to using a fixed set of exemplars for all test queries submitted to the LLM provider (e.g., via an API), using a different set of exemplars for every test query may lead to greater data exposure, i.e., more data exemplars being exposed to the LLM provider. Privacy issues have become increasingly prominent in discussions surrounding LLMs [22], with evidence of user data being output by these models [36]. Therefore, retrieval-based methods are undesirable in scenarios where such privacy concerns are important considerations. Given these two important shortcomings of retrieval-based methods, it is imperative to develop exemplar selection methods that generalize to test queries and are capable of choosing a fixed set of exemplars that perform well for all test queries for a task.
@@
In this work, we propose a novel exemplar selection algorithm that addresses the above-mentioned challenges faced by existing works (with detailed discussions of related work in Sec. 2). We formulate exemplar selection as a black-box optimization problem, in which every input in the domain corresponds to a sequence of $k$ exemplars and its corresponding output represents the ICL performance achieved by including this sequence in the prompt for the LLM. Given this formulation, for every sequence of $k$ exemplars in the domain, we adopt the embedding from a powerful pre-trained language model as its continuous representation, and train a neural network (NN) to predict its ICL performance. Based on the trained NN, we use a neural bandit algorithm to find the optimal exemplar sequence in a query-efficient manner. Our Efficient ordering-aware Automated Selection of Exemplars (EASE) algorithm offers several significant benefits:
@@
2 Related work
Instruction optimization. Another related direction for enhancing the performance of LLMs is instruction optimization. Focusing on instructions within the prompt, evolutionary algorithms and zeroth-order optimization algorithms are gaining popularity in refining the prompts for black-box LLMs [4, 14, 8, 6, 37, 10]. Specifically, some studies [14, 6, 37, 10] maintained a constant set of exemplars throughout the process of prompt optimization, whereas others [4, 8] ignored the consideration of exemplars altogether by adopting a zero-shot setting. It is therefore imperative to develop an effective exemplar selection method for black-box LLMs.
@@ 4.1 NeuralUCB for query-efficient optimization of exemplar sequences
Inspired by Lin et al. [14] who have shown impressive performances of the neural upper confidence bound (NeuralUCB) acquisition function [42] on instruction optimization, we adopt the NeuralUCB approach to our novel black-box exemplar selection formulation in (1). We propose to use a neural network (NN) to directly learn the mapping from a general-purpose hidden embedding of input exemplar sequences to the validation score. Specifically, we propose to use a embedding model $h(\cdot)$ and train a network $m(h(E);\theta_{t})$ at iteration $t$ such that
@@ 4.3 Natural extension to jointly optimize instructions and exemplars
Our algorithm can further boost the performance of ICL for LLMs by jointly optimizing the exemplars and the instruction. A natural extension of our formulation in Sec. 3 allows the instruction, being another essential component of the LLM prompt, to be simultaneously optimized with exemplars. This ensures the optimal matching between instructions and exemplars to achieve a fully automated pipeline with superior performance. Specifically, our formulation naturally extends to $E=(p,e_{1},e_{2},\ldots,e_{k})$ where $p\in P$ is an instruction from a candidate space/set $P$, i.e., $Q_{t}^{\prime}\leftarrow P\times Q_{t}^{\prime}$. Subsequently, $p$ can be intuitively treated as another type of “exemplar” from a new space $P$ and optimized in conjuction with the exemplars. In practice, the fixed set $P$ of candidate instructions can be generated by the black-box model through techniques such as APE [43], PromptBreeder [6], etc. This extension is not only simple in its implementation but also proven to be effective in our experiments in Sec. 5.4.
@@ 5 Experiments
We compare EASE with the following comprehensive suite of baselines. They include subset selection methods with (a) a determinantal point process (DPP) metric adapted from [38], and (b) maximum mean discrepancy (MMD[30], (c) optimal transport (OT[33] metrics. We also adapt retrieval-based methods to our setting using a new retrieve-then-sample strategy based on the validation set (details in App. B.2 and Sec. 5.5), specifically with the classical (d) Cosine similarity and (e) BM25 [28] retrievers. We also compare with existing exemplar selection baselines using (f) an active selection policy learned using reinforcement learning (Active[39], and (g) an exemplar influence metric (Inf[23]. Additionally, we propose two more new strong baselines: (h) Evo which mutates exemplars through evolutionary strategies and (i) Best-of-N which explores the exemplar space uniformly until the query budget is exhausted. More implementation details are found in App. B.2. The number of exemplars in the in-context prompt is set to $k=5$. The black-box query budget is 165 evaluations following Lin et al. [14]. Since the effectiveness of the optimization is directly reflected by the value of the objective function in (1), we report the validation accuracy in the following experiments unless otherwise specified. The test accuracy tables are presented in App. C.10.
@@ 5.1 Empirical study of performance gains on various benchmark tasks
To study the effectiveness of EASE, we conduct empirical comparisons with baseline methods using the Instruction Induction (II) benchmark tasks [4]. We test on tasks that contain more than 100 training examples for selection. The results are shown in Tab. 1. Our EASE achieves the best performance in 17 out of 19 language tasks. We observe that the subset selection methods such as DPP, MMD and OT are ineffective in practice, implying that choosing subsets alone without considering the order information is inadequate. While the retrieval-based methods Cosine and BM25 have demonstrated success in retrieving test-sample-based exemplars [29], it is not as effective in finding the best exemplars that generalize to the entire task. This is because the exemplars retrieved at the instance level may not work well for the entire validation set. The Active baseline requires training an active exemplar selection policy through the data-intensive reinforcement learning process, which explains its ineffectiveness under our budget-constrained setting. The Inf baseline is not efficient in subset sampling and disregards exemplar ordering, leading to worse results. Our proposed Evo and Best-of-N are the most competitive baselines with best performance in 9 and 11 tasks, respectively.
@@
antonyms 70.0$\pm$0.0 80.0$\pm$0.0 81.7$\pm$1.4 85.0$\pm$0.0 85.0$\pm$0.0 80.0$\pm$0.0 86.7$\pm$1.4 88.3$\pm$1.4 90.0$\pm$0.0 90.0$\pm$0.0
auto_categorization 3.3$\pm$1.4 8.3$\pm$1.4 0.0$\pm$0.0 25.0$\pm$0.0 16.7$\pm$1.4 10.0$\pm$2.4 21.7$\pm$1.4 21.7$\pm$1.4 20.0$\pm$0.0 30.0$\pm$0.0
@@
object_counting 55.0$\pm$2.4 56.7$\pm$1.4 48.3$\pm$1.4 61.7$\pm$1.4 66.7$\pm$1.4 51.7$\pm$1.4 63.3$\pm$3.6 70.0$\pm$0.0 70.0$\pm$0.0 73.3$\pm$1.4
orthography_starts_with 20.0$\pm$2.4 35.0$\pm$0.0 61.7$\pm$1.4 78.3$\pm$1.4 70.0$\pm$0.0 43.3$\pm$1.4 70.0$\pm$2.4 75.0$\pm$0.0 78.3$\pm$1.4 78.3$\pm$1.4
rhymes 60.0$\pm$0.0 51.7$\pm$1.4 0.0$\pm$0.0 100.0$\pm$0.0 80.0$\pm$0.0 65.0$\pm$8.2 70.0$\pm$10.8 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
second_word_letter 10.0$\pm$2.4 30.0$\pm$0.0 28.3$\pm$1.4 50.0$\pm$0.0 50.0$\pm$0.0 26.7$\pm$8.3 40.0$\pm$0.0 46.7$\pm$1.4 50.0$\pm$0.0 50.0$\pm$0.0
sentence_similarity 20.0$\pm$0.0 21.7$\pm$2.7 40.0$\pm$2.4 46.7$\pm$1.4 53.3$\pm$1.4 5.0$\pm$4.1 18.3$\pm$5.4 45.0$\pm$0.0 51.7$\pm$1.4 56.7$\pm$1.4
sentiment 85.0$\pm$0.0 90.0$\pm$0.0 85.0$\pm$0.0 96.7$\pm$1.4 100.0$\pm$0.0 85.0$\pm$4.1 91.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
@@
synonyms 10.0$\pm$0.0 25.0$\pm$0.0 20.0$\pm$0.0 35.0$\pm$0.0 30.0$\pm$0.0 3.3$\pm$1.4 26.7$\pm$1.4 30.0$\pm$0.0 30.0$\pm$0.0 30.0$\pm$0.0
taxonomy_animal 43.3$\pm$3.6 40.0$\pm$2.4 46.7$\pm$1.4 85.0$\pm$2.4 80.0$\pm$0.0 45.0$\pm$6.2 70.0$\pm$4.1 80.0$\pm$0.0 80.0$\pm$0.0 88.3$\pm$2.7
@@
translation_en-es 90.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.4 98.3$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
translation_en-fr 76.7$\pm$1.4 76.7$\pm$1.4 81.7$\pm$1.4 85.0$\pm$0.0 85.0$\pm$0.0 81.7$\pm$1.4 85.0$\pm$0.0 86.7$\pm$1.4 85.0$\pm$0.0 88.3$\pm$1.4
word_sorting 26.7$\pm$1.4 88.3$\pm$1.4 88.3$\pm$1.4 90.0$\pm$0.0 71.7$\pm$1.4 80.0$\pm$0.0 88.3$\pm$1.4 93.3$\pm$1.4 91.7$\pm$1.4 91.7$\pm$1.4
word_unscrambling 68.3$\pm$1.4 56.7$\pm$1.4 71.7$\pm$1.4 75.0$\pm$0.0 76.7$\pm$1.4 63.3$\pm$3.6 66.7$\pm$1.4 75.0$\pm$0.0 75.0$\pm$0.0 78.3$\pm$2.7
# best-performing tasks 2 2 2 8 5 1 5 9 11 17

### v2

@@ Abstract
Large language models (LLMs) have shown impressive capabilities in real-world applications. The capability of in-context learning (ICL) allows us to adapt an LLM to downstream tasks by including input-label exemplars in the prompt without model fine-tuning. However, the quality of these exemplars in the prompt greatly impacts performance, highlighting the need for an effective automated exemplar selection method. Recent studies have explored retrieval-based approaches to select exemplars tailored to individual test queries, which can be undesirable due to extra test-time computation and an increased risk of data exposure. Moreover, existing methods fail to adequately account for the impact of exemplar ordering on the performance. On the other hand, the impact of the instruction, another essential component in the prompt given to the LLM, is often overlooked in existing exemplar selection methods. To address these challenges, we propose a novel method named EASE, which leverages the hidden embedding from a pre-trained language model to represent ordered sets of exemplars and uses a neural bandit algorithm to optimize the sets of exemplars while accounting for exemplar ordering. Our EASE can efficiently find an ordered set of exemplars that performs well for all test queries from a given task, thereby eliminating test-time computation. Importantly, EASE can be readily extended to jointly optimize both the exemplars and the instruction. Through extensive empirical evaluations (including novel tasks), we demonstrate the superiority of EASE over existing methods, and reveal practical insights about the impact of exemplar selection on ICL performance, which may be of independent interest. Our code is available at https://github.com/ZhaoxuanWu/EASE-Prompt-Optimization.
@@ 1 Introduction
00footnotetext: * Equal contribution. Correspondence to: Zhongxiang Dai <daizhongxiang@cuhk.edu.cn>.
Large language models (LLMs) have recently drawn significant attention and have been widely deployed in various real-world scenarios [26, 28, 51] due to their strong capabilities. Of note, a particularly impressive capability of LLMs is in-context learning (ICL): LLMs can learn from a handful of input-label demonstrations (i.e., data exemplars) included in its prompt to perform a downstream task [3, 23]. ICL allows us to adapt an LLM to a downstream task without fine-tuning the model parameters [20, 29]. However, the ICL performance is heavily dependent on the data exemplars in the prompt [1, 23, 38]. Therefore, to maximize the ICL performance, it is of paramount importance to carefully select the set of exemplars [46]. Unfortunately, exemplar selection for ICL is challenging because the mechanism of ICL is complicated and unknown, and this difficulty is further aggravated for black-box LLMs to which we only have API access (e.g., GPT-4 [32], Gemini Ultra [41]).
A large number of existing works on exemplar selection for ICL have considered retrieval-based methods [1, 23, 38]. Specifically, they aim to develop a retriever model such that for every test query, the retriever model retrieves the corresponding best set of exemplars tailored to this particular query [1, 48]. Such approaches require test-time retrieval for every query which might hinder their ease of adoption [25] while test-time computation is not needed for our method. Moreover, another drawback of these retrieval-based methods is that they may lead to increased privacy risks. Compared to using a fixed set of exemplars for all test queries submitted to the LLM provider (e.g., via an API), using a different set of exemplars for every test query may lead to greater data exposure, i.e., more data exemplars being exposed to the LLM provider. Privacy issues have become increasingly prominent in discussions surrounding LLMs [30], with evidence of user data being output by these models [45]. Therefore, retrieval-based methods are undesirable in scenarios where such privacy concerns are important considerations. Given these two important shortcomings of retrieval-based methods, it is imperative to develop exemplar selection methods that generalize to test queries and are capable of choosing a fixed set of exemplars that perform well for all test queries for a task.
@@
In this work, we propose a novel exemplar selection algorithm that addresses the above-mentioned challenges faced by existing works (with detailed discussions of related work in App. 5). We formulate exemplar selection as a black-box optimization problem, in which every input in the domain corresponds to a sequence of $k$ exemplars and its corresponding output represents the ICL performance achieved by including this sequence in the prompt for the LLM. Given this formulation, for every sequence of $k$ exemplars in the domain, we adopt the embedding from a powerful pre-trained language model as its continuous representation, and train a neural network (NN) to predict its ICL performance. Based on the trained NN, we use a neural bandit algorithm to find the optimal exemplar sequence in a query-efficient manner. Our Efficient ordering-aware Automated Selection of Exemplars (EASE) algorithm offers several significant benefits:
@@ 3.1 NeuralUCB for query-efficient optimization of exemplar sequences
Inspired by Lin et al. [21] who have shown impressive performances of the neural upper confidence bound (NeuralUCB) acquisition function [53] on instruction optimization, we adopt the NeuralUCB approach to our novel black-box exemplar selection formulation in (1). We propose to use a neural network (NN) to directly learn the mapping from a general-purpose hidden embedding of input exemplar sequences to the validation score. Specifically, we propose to use an embedding model $h(\cdot)$ and train a network $m(h(E);\theta_{t})$ at iteration $t$ such that
@@ 3.3 Natural extension to jointly optimize instructions and exemplars
Our algorithm can further boost the performance of ICL for LLMs by jointly optimizing the exemplars and the instruction. A natural extension of our formulation in Sec. 2 allows the instruction, being another essential component of the LLM prompt, to be simultaneously optimized with exemplars. This ensures the optimal matching between instructions and exemplars to achieve a fully automated pipeline with superior performance. Specifically, our formulation naturally extends to $E=(p,e_{1},e_{2},\ldots,e_{k})$ where $p\in P$ is an instruction from a candidate space/set $P$, i.e., $Q_{t}^{\prime}\leftarrow P\times Q_{t}^{\prime}$. Subsequently, $p$ can be intuitively treated as another type of “exemplar” from a new space $P$ and optimized in conjunction with the exemplars. In practice, the fixed set $P$ of candidate instructions can be generated by the black-box model through techniques such as APE [54], PromptBreeder [10], etc. This extension is not only simple in its implementation but also proven to be effective in our experiments in Sec. 4.4.
@@ 4 Experiments
We compare EASE with the following comprehensive suite of baselines. They include subset selection methods with (a) a determinantal point process (DPP) metric adapted from [48], and (b) maximum mean discrepancy (MMD[39], (c) optimal transport (OT[42] metrics. We also adapt retrieval-based methods to our setting using a new retrieve-then-sample strategy based on the validation set (details in App. B.2 and Sec. 4.5), specifically with the classical (d) Cosine similarity and (e) BM25 [37] retrievers. We also compare with existing exemplar selection baselines using (f) an active selection policy learned using reinforcement learning (Active[50], and (g) an exemplar influence metric (Inf[31]. Additionally, we propose two more new strong baselines: (h) Evo which mutates exemplars through evolutionary strategies and (i) Best-of-N which explores the exemplar space uniformly until the query budget is exhausted. Evo is similar to PhaseEvo proposed by Cui et al. [8]. Best-of-N shares similarity with DSPy’s BootstrapFewShot with RandomSearch [17]. More implementation details are found in App. B.2. The number of exemplars in the in-context prompt is set to $k=5$. The black-box query budget is 165 evaluations following Lin et al. [21]. Since the effectiveness of the optimization is directly reflected by the value of the objective function in (1), we report the validation accuracy in the following experiments unless otherwise specified. The test accuracy tables are presented in App. C.15.
@@ 4.1 Empirical study of performance gains on various benchmark tasks
To study the effectiveness of EASE, we conduct empirical comparisons with baseline methods using the Instruction Induction (II) benchmark tasks [5]. We test on tasks that contain more than 100 training examples for selection. The results are shown in Tab. 1. Our EASE achieves the best performance in 17 out of 19 language tasks. We observe that the subset selection methods such as DPP, MMD and OT are ineffective in practice, implying that choosing subsets alone without considering the order information is inadequate. While the retrieval-based methods Cosine and BM25 have demonstrated success in retrieving test-sample-based exemplars [38], it is not as effective in finding the best exemplars that generalize to the entire task. This is because the exemplars retrieved at the instance level may not work well for the entire validation set. The Active baseline requires training an active exemplar selection policy through the data-intensive reinforcement learning process, which explains its ineffectiveness under our budget-constrained setting. The Inf baseline is not efficient in subset sampling and disregards exemplar ordering, leading to worse results. Our proposed Evo and Best-of-N are the most competitive baselines both with best performance in 9 tasks.
@@
antonyms 70.0$\pm$0.0 80.0$\pm$0.0 81.7$\pm$1.7 85.0$\pm$0.0 85.0$\pm$0.0 80.0$\pm$0.0 86.7$\pm$1.7 88.3$\pm$1.7 90.0$\pm$0.0 90.0$\pm$0.0
auto_categorization 3.3$\pm$1.7 8.3$\pm$1.7 0.0$\pm$0.0 25.0$\pm$0.0 16.7$\pm$1.7 10.0$\pm$2.4 21.7$\pm$1.7 21.7$\pm$1.7 20.0$\pm$0.0 30.0$\pm$0.0
@@
object_counting 55.0$\pm$2.9 56.7$\pm$1.7 48.3$\pm$1.7 61.7$\pm$1.7 66.7$\pm$1.7 51.7$\pm$1.4 63.3$\pm$4.4 70.0$\pm$0.0 70.0$\pm$0.0 73.3$\pm$1.7
orthography_starts_with 20.0$\pm$2.9 35.0$\pm$0.0 61.7$\pm$1.7 78.3$\pm$1.7 70.0$\pm$0.0 43.3$\pm$1.4 70.0$\pm$2.9 75.0$\pm$0.0 78.3$\pm$1.7 80.0$\pm$0.0
rhymes 60.0$\pm$0.0 51.7$\pm$1.7 0.0$\pm$0.0 100.0$\pm$0.0 80.0$\pm$0.0 65.0$\pm$8.2 70.0$\pm$13.2 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
second_word_letter 10.0$\pm$2.9 30.0$\pm$0.0 28.3$\pm$1.7 50.0$\pm$0.0 50.0$\pm$0.0 26.7$\pm$8.3 40.0$\pm$0.0 46.7$\pm$1.7 50.0$\pm$0.0 53.3$\pm$1.7
sentence_similarity 20.0$\pm$0.0 21.7$\pm$3.3 40.0$\pm$2.9 46.7$\pm$1.7 53.3$\pm$1.7 5.0$\pm$4.1 18.3$\pm$6.7 45.0$\pm$0.0 51.7$\pm$1.7 56.7$\pm$1.7
sentiment 85.0$\pm$0.0 90.0$\pm$0.0 85.0$\pm$0.0 96.7$\pm$1.7 100.0$\pm$0.0 85.0$\pm$4.1 91.7$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
@@
synonyms 10.0$\pm$0.0 25.0$\pm$0.0 20.0$\pm$0.0 35.0$\pm$0.0 30.0$\pm$0.0 3.3$\pm$1.4 26.7$\pm$1.7 30.0$\pm$0.0 30.0$\pm$0.0 30.0$\pm$0.0
taxonomy_animal 43.3$\pm$4.4 40.0$\pm$2.9 46.7$\pm$1.7 85.0$\pm$2.9 80.0$\pm$0.0 45.0$\pm$6.2 70.0$\pm$5.0 80.0$\pm$0.0 80.0$\pm$0.0 88.3$\pm$1.7
@@
translation_en-es 90.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.4 98.3$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
translation_en-fr 76.7$\pm$1.7 76.7$\pm$1.7 81.7$\pm$1.7 85.0$\pm$0.0 85.0$\pm$0.0 81.7$\pm$1.4 85.0$\pm$0.0 86.7$\pm$1.7 85.0$\pm$0.0 88.3$\pm$1.7
word_sorting 26.7$\pm$1.7 88.3$\pm$1.7 88.3$\pm$1.7 90.0$\pm$0.0 71.7$\pm$1.7 80.0$\pm$0.0 88.3$\pm$1.7 93.3$\pm$1.7 91.7$\pm$1.7 90.0$\pm$0.0
word_unscrambling 68.3$\pm$1.7 56.7$\pm$1.7 71.7$\pm$1.7 75.0$\pm$0.0 76.7$\pm$1.7 63.3$\pm$3.6 66.7$\pm$1.7 75.0$\pm$0.0 75.0$\pm$0.0 78.3$\pm$1.7
# best-performing tasks 2 2 2 6 4 1 5 9 9 17